OpenStudy (anonymous):

EMERGENCY! :( Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences. Btw, f(x)= x^2+6x - 16

OpenStudy (mathmate):

Let \(f(x)=x^2+6x-16=ax^2+bx+c\) be the standard form, then a=1, b=6, c=-16 the equation in vertex form is \(\large f(x)=a(x-h)^2+k\) where a has the same value in both forms, and where \(\large h=-\frac{b}{2a},~and~k=c-ah^2\) Substitute a, b and c in the preceding equations of h and k to find the vertex form. If you would like to know how to do the conversion without memorizing the above formulas, use completing the square, or simply equate the two forms and solve for a, h and k.